OpenStudy (anonymous):

Which expressions are equivalent to the one below? Check all that apply. log(10^5) A. 1 B. 5 C. 5*10 D. 5*log10

OpenStudy (mathmale):

This problem might actually be easier if expressed as \[\log 10^5\]

OpenStudy (mathmale):

there is a rule of logs that goes like this:\[\log a^b = b*\log a\] See what you can do with this info.